Use Karnaugh map to minimize the sum of product expansion
xy'z+ xy'z'+x'yz+x'y'z+x'y'z'
1
Expert's answer
2020-10-25T18:56:30-0400

"\text{create a Karnaugh map}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \frac{yz}{x} & 00 &01&11&10 \\ \hline
 0 & 1 & 1&1&0 \\
 \hdashline
 1 & 1& 1&0&0
\end{array}"

"\text{Let's select on the Karnaugh map rectangular areas of units of the largest area,}"

"\text{which are powers of two, and write out the conjunctions corresponding to them:}"


"\text{region 1}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \frac{yz}{x} & 00 &01&11&10 \\ \hline
 0 & \color{red}1 &\color{red} 1&1&0 \\
 \hdashline
 1 & \color{red}1& \color{red}1&0&0
\end{array}"

"K_1:y'"


"\text{region 2}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \frac{yz}{x} & 00 &01&11&10 \\ \hline
 0 & 1 & \color{red}1&\color{red}1&0 \\
 \hdashline
 1 & 1& 1&0&0
\end{array}"

"K_2:x^{\prime}z"

"\text{Combining them using the OR operation, we get}"

"y'+x'z"

Answer: "y'+x'z"